# Neural machine translation of clinical procedure codes for medical diagnosis and uncertainty quantification

### v1

@@ Abstract
Index Terms: Clinical Decision Support System (CDSS), uncertainty quantification (UQ), medical entropy, entropy reduction, diagnosis predictor
@@ 2 Materials and Methods
2.1 Data Source
@@ 2.4 Seq2seq diagnoses predictor
A predictive model comes in handy for predicting diagnoses based on procedures that a patient received in an admission. Notably, both diagnosis and procedure codes are in sequential form. Namely, procedure codes possess information in chronological order, whereas the order in diagnosis codes reveals the importance of the patient’s condition. Consequently, we adopt the seq2seq model to obtain a more robust prediction of diagnoses.
The seq2seq model [15]is a framework architecture that accepts a source sequence as input and generates a target sequence as output, composed of an encoder and a decoder. Instead of being confined by the fixed length of input and output sequences, the seq2seq framework enables us to employ flexible sequence lengths, which makes it robust for the scenario between diagnoses and procedures. Specifically, the encoder takes the input sequence to produce a “context”, an embedding vector composed of contextual information of the source sequence. Afterward, the decoder outputs the target sequence based on the context.
Here, both the decoder and encoder are recurrent neural networks (RNNs). For every input procedure, the encoder of the seq2seq diagnoses predictor outputs a hidden state, then applies the hidden state to the following input procedure. When the procedure ends, the final procedure’s hidden state that encodes the entire procedure sequence is precisely the context vector. At the beginning of decoding, the decoder takes the context vector from the encoder as the initial hidden state. During every decoding step, the decoder outputs a most-likely diagnosis based on the previously predicted diagnosis and the current hidden state.
To acquire the distribution of probable diagnoses depending on the different stages during admission for the downstream entropy task, we are more concerned with the distribution of the probable diagnosis (i.e., in terms of the ICD-9 codes). During training, the procedure ICD-9 codes are fed as input to the encoder, while the diagnostic ICD-9 codes are the output of the decoder. For evaluation, we could apply the testing procedure and diagnose ICD codes to evaluate the model performance and use the decoder’s softmax output, which provides us with a more informative tool by quantifying information entropy.
2.4.1 Attention mechanism
Instead of receiving the last hidden state of the encoder as input, the decoder with an attention mechanism takes all the hidden states as input[20, 21]. Before generating output, the decoder scores different hidden states based on correlation and adds them by their corresponding weights. Thus, the decoder could predict the target based on the weighted-summed hidden states, thereby paying more attention to the hidden states that are of higher importance for an accurate prediction of the target.
2.4.2 Teacher Forcing
The prediction capability during the early stage of RNN training is insufficient to generate good results. Moreover, this issue causes error propagation and makes model training too slow to converge. Instead of using the previous output for the following state input, we adopt the teacher-forcing mechanism[22], which directly adopts the ground truth as input for training.
2.4.3 Entropy Quantification in Admissions
@@ 3 Results
Table 2: Entropy of most frequent first N=1 procedure.
@@ 4 Discussion
Table 3: Entropy of most frequent first N=2 procedures
@@
Table 4: Entropy of most frequent first N=3 procedures
The findings from Tables 2,  3, and  4 bring additional insights into how medical procedures affect diagnostic uncertainty. For instance, Table 2 points out the inherent differences in uncertainty between procedures like "7569" (related to baby delivery) and "9671" (used for patients with respiratory failure). The former typically has more straightforward diagnoses compared to the latter, which is affected by various unknown factors.
Table 3 shows the entropy drops in similar scenarios but for the first two procedures from the initial admission. There are interesting and informative findings. First, when we focus on the first and the fifth combination, they start with the same procedure, "0066" known as “percutaneous transluminal coronary angioplasty (PTCA)” and receive roughly the same following procedure, "3607" and "3606", which stand for "Insertion of drug-eluting coronary artery stent(s)" and "Insertion of non-drug-eluting coronary artery stent(s)". Notably, the entropy when receiving the second procedure shows that the proposed framework is capable of differentiating the two procedure combinations. On the other hand, we observe that the entropy of the "3722-8856" combination decreases from 10.37 to 4.55, which is the largest difference among the top ten combinations, where "3722" stands for "Left heart cardiac catheterization" and "8856" for "Coronary arteriography using two catheters." Compared with other procedures, coronary arteriography is an invasive procedure with more in-depth informative outcomes, showing imaging of blocked blood vessels causing coronary artery disease. Thus, this might be reasonable why coronary arteriography followed by cardiac catheterization could scale down the uncertainties to this level.
Table 4 extends this observation to the first three procedures received by a patient. These ten trends start from the same initial entropy and then act nearly the same when receiving the first procedure but act entirely differently after receiving three procedures. When the patient receives the third procedure, the entropy of the procedure triplet "3612-3615-3961", which involves critical coronary bypass surgeries and auxiliary circulation for open heart surgery, drops dramatically to 1.4552; in contrast, the entropy of triplet "8952-8744-8938" which are “Electrocardiogram”, “Nonoperative respiratory measurement” and “Routine chest X-ray” declines slightly to 6.9336. Similarly, the nature of these three triplets can explain this. To specify, although the two triplets are related to heart disease, the triplet "3612-3615-3961" is a set of imperative procedures in life-threatening situations, and the triplet "8952-8744-8938" is used for preliminary examination for heart-related disease.
The entropy trends in Figure 4, the three cases suggest varied impacts of medical procedures on patient outcomes. The initial entropy reduction in Admission #1 indicates the effectiveness of the first procedure in reducing uncertainty, but subsequent procedures did not contribute significantly to further reducing the entropy, leading to a constant level until the patient’s demise. This may reflect the severity of the patient’s condition and the limited impact of later treatments.
Admission #2’s entropy trend, with its dramatic initial drop and subsequent fluctuations, suggests that the sequence and nature of procedures can significantly impact the entropy. The fluctuating pattern indicates varying degrees of effectiveness of the procedures in reducing uncertainty. The final decrease in entropy suggests that the later procedures were more informative or effective in clarifying the patient’s condition. In contrast, the steady decrease in entropy in Admission #3 implies a consistent reduction in uncertainty with each procedure, possibly indicating a more straightforward or predictable case. This pattern suggests that in certain cases, a series of procedures can progressively contribute to a better understanding of the patient’s condition, aiding in decision-making and prognosis.
These observations underscore the importance of procedure sequencing and selection in managing complex cases like sepsis, where comorbidities and unpredictable outcomes are common. Understanding these entropy trends can help physicians make more informed decisions, moving away from a "black box" approach to a more transparent and explainable medical practice.
The proposed framework’s application in both simulated and real-world scenarios suggests its potential usefulness across different medical cases. Using ICD codes for data analysis helps address issues like missing values and outliers, although this approach primarily focuses on standardization and reliability. The study’s findings have implications for the potential enhancement of CDSS through the incorporation of data-driven AI/ML models. By offering a measure of medical entropy, the framework might assist clinicians in making effective and robust decisions during a patient’s hospital stay, especially in settings where time is critical and the uncertainties may be significant.
However, we note that our current study is not without its limitations. One of the key challenges faced was the potential sub-optimality of the seq2seq model in certain clinical scenarios. While effective in our study, this model may not be universally applicable in all settings, suggesting the need for further research into alternative predictive models. Additionally, the reliance on ICD codes, while beneficial in some respects, may limit the depth and breadth of insights compared to analyses that incorporate more detailed patient data. In this study, ICD-9 and CPT codes have limitations as they are generated post patient stay. Therefore, their use in our framework is primarily for validation purposes only.
@@
For future studies, we propose utilizing Electronic Health Records (EHR) audit logs for a more discrete measure of entropy quantification and reduction throughout hospital visits. This approach could provide a more nuanced understanding of the patient’s journey, capturing the intricacies of clinical decision-making in real-time. EHR audit logs, with their detailed and sequential data, offer a rich resource for comprehensively analyzing the entropy changes during a patient’s hospital stay, potentially enhancing the predictive power and applicability of our framework in various clinical environments[27, 28, 29].
@@
In summary, this study offers a novel perspective on using a computational framework to approach medical uncertainty in clinical settings. While the framework shows promise in adapting to various clinical scenarios, its broader impact on healthcare and patient care, particularly in the development of Clinical Decision Support Systems, remains an area for further exploration.
@@ 5 Conclusion
6 Author contributions
PC, SH, BY proposed the concept, developed the framework, and designed the study. PC implemented the model, performed the experiments, and drafted the initial manuscript. PC, SH, BY, NK, AA analyzed the results and revised the manuscript. SH, BY supervised the overall study.
7 Competing interests
The authors have no competing interests to declare.
Appendix A In-depth analysis for the three sepsis cases
Figure 5: Distribution changes when receiving procedures during hospitalization. H stands for entropy and the procedure sequences are for admission 1 shown in Figure 3
Figure 6: Distribution changes when receiving procedures during hospitalization. H stands for entropy and the procedure sequences are for admission 2 shown in Figure 3
Figure 7: Distribution changes when receiving procedures during hospitalization. H stands for entropy and the procedure sequences are for admission 3 shown in Figure 3

### v2

@@ 2 Materials and Methods
2.1 Data Source and Experiment Compute Resource
@@
All experiments in this study were conducted on a CPU, using an Apple MacBook Pro equipped with the M1 Max chip. This setup not only provided sufficient computational power for processing the ICD-9 coded data and running the machine learning models, but also enabled us to carry out the experiments with relative ease, without requiring additional high-performance computing resources such as GPUs or cloud-based servers. This also ensured that the study remains reproducible on accessible hardware for users with similar computational setups.
@@ 2.4 Seq2seq diagnoses predictor
To predict diagnoses based on procedures during patient admission, we use a Seq2Seq model. Both diagnosis and procedure codes are sequential, with procedure codes following a chronological order, while the order of diagnosis codes reflects the severity of the patient’s condition. The Seq2Seq model offers a flexible approach to capture this dynamic by processing variable-length sequences, making it ideal for linking procedures to diagnoses.
The Seq2Seq framework, based on the encoder-decoder architecture [15], takes a source sequence as input and generates a target sequence as output. The encoder processes the input sequence to create a context vector, an embedding of the entire source sequence, which is then passed to the decoder to generate the target sequence. In our case, the encoder processes the procedure sequence, and its final hidden state serves as the context vector. The decoder uses this context to predict diagnoses sequentially, with each prediction informing the next.
We also integrate an attention mechanism [20, 21] to allow the decoder to weigh and focus on the most relevant hidden states from the encoder, enhancing prediction accuracy. To further improve training efficiency and prevent error propagation, we apply the teacher-forcing technique [22], which inputs the ground truth directly during training.
We focus on predicting the distribution of diagnoses at different stages of admission using ICD-9 codes. The encoder is fed with procedure ICD-9 codes, while the decoder outputs diagnosis ICD-9 codes. During evaluation, the softmax output of the decoder provides a probabilistic distribution of diagnoses, aiding in entropy-based analyses.
2.5 Entropy Quantification in Admissions
@@ 3 Results
Table 2: Entropy of most frequent first N=1 procedure.
@@ 4 Discussion
Table 3: Entropy of most frequent first N=2 procedures
@@
Table 4: Entropy of most frequent first N=3 procedures
@@
The findings from Tables 2,  3, and 4 provide insights into how medical procedures influence diagnostic uncertainty. For instance, Table2 highlights differences in uncertainty between procedures like "7569" (related to delivery) and "9671" (used for respiratory failure). The former typically leads to more straightforward diagnoses, while the latter involves more uncertainty due to various unknown factors. Table 3 examines entropy drops for the first two procedures after admission. Interestingly, procedures "0066" (PTCA) paired with either "3607" (drug-eluting stent) or "3606" (non-drug-eluting stent) show that the framework can differentiate between these combinations. Additionally, the "3722-8856" pair ("Left heart catheterization" and "Coronary arteriography") shows the largest entropy drop, reflecting the detailed information provided by coronary arteriography. Table 4 extends this analysis to the first three procedures. The entropy for triplet "3612-3615-3961" (critical coronary bypass surgeries) drops significantly, while triplet "8952-8744-8938" (preliminary heart-related exams) shows only a modest decline. This contrast underscores the importance of life-saving procedures in reducing uncertainty. The entropy trends in Figure 4 demonstrate varied impacts of procedures on patient outcomes. In Admission #1, initial entropy reduction reflects the effectiveness of the first procedure, but later procedures contribute little, suggesting the severity of the condition. Admission #2 shows fluctuating entropy, indicating varying effectiveness of the procedures, with a final drop suggesting more informative interventions. Admission #3 displays a steady decrease in entropy, indicating consistent reduction in uncertainty with each procedure.
These findings highlight the importance of procedure sequencing and selection in managing complex cases like sepsis, where uncertainty is high. Understanding entropy trends can help physicians make more informed decisions, offering a more transparent and explainable approach to medical practice. The proposed framework, tested in both simulated and real-world scenarios, demonstrates its potential usefulness across diverse medical cases. The use of ICD codes addresses issues like missing values and outliers, promoting standardization and reliability. These insights have implications for enhancing Clinical Decision Support Systems (CDSS) by integrating data-driven AI/ML models. By quantifying medical entropy, the framework may assist clinicians in making robust, timely decisions, particularly in high-uncertainty situations. However, this study has some limitations. The seq2seq model, while effective here, may not be optimal in all clinical scenarios, and further exploration of alternative models is necessary. Additionally, relying on ICD codes may limit the depth of analysis, as they are generated after patient stays. Their use here is primarily for validation.
@@
Future research could leverage Electronic Health Records (EHR) audit logs to provide a more granular measure of entropy throughout hospital visits, capturing real-time decision-making processes. EHR audit logs offer rich, sequential data that could enhance the predictive power and applicability of our framework in clinical environments [27, 28, 29].
@@
In summary, this study introduces a novel framework for addressing medical uncertainty in clinical settings. While the approach shows promise in various scenarios, further research is needed to explore its broader impact on healthcare, particularly in advancing Clinical Decision Support Systems.